OpenStudy (anonymous):

r(radius)=__ d(diameter)=__ c(circumference)=19 in. a(area)=__ how do you find out the radius and diameter and area?

OpenStudy (anonymous):

c=2pi *r a=pi r^2

OpenStudy (anonymous):

d=2r

OpenStudy (lgbasallote):

c = pi d d = c/pi d = 19/3.14.16 d = 6.05 r = 3.025 area = pi r^2 area = pi (3.025)^2 = pi (9.15) area = 28.75